# Specification and Evaluation of Multi-Agent LLM Systems - Prototype and Cybersecurity Applications

### v3

@@ IV-B2 Reasoning Chains
can be constructed based on a series of prompts and responses together with constraints [16, 14]. In addition to Chain-of-Thought (CoT), which aims to issue prompts that generate subsequent reasoning steps with their responses, the chain can also be constructed explicitly by relying on pre-defined structures. Constraints ensure their existence and allow prompting the LLM of an agent with the next step. This is especially relevant in multi-agent scenarios, where another, specialized agent might be invoked based on previously collected data. For example, Listing 2 shows the Server-Security-Agent type that establishes an IP address, network, and scan result data after a series of prompts and invokes an agent of type Audit-Report-Agent.
On the basis of these prompts, the agent is prompted to create a report using the passed data. Further actions might place constraints or specific conditions influencing which agent is invoked or the data passed to the agent. Figure 3 in Appendix Specification and Evaluation of Multi-Agent LLM Systems - Prototype and Cybersecurity Applications shows an excerpt of the generated report.
@@ V Conclusion
This paper presented the initial results of experimental research on the specification of multi-agent LLM systems. In this first system overview, the architecture and specification language were demonstrated and applied in test cases related to question answering as well as network and server security tasks. The test cases indicate feasibility of the architecture and specification, showing the completion of tasks with agent specifications based on state-of-the-art commercially and openly available LLMs. In particular, the completion of the test cases shows the potential of LLM agents for software-based tasks in applications. Factors enabling the task completion are a combination of (1.) involving the knowledge capabilities of specialized agents such as for cybersecurity and code generation, (2.) utilizing reasoning techniques such as reasoning chains, and (3.) agents executing the inferred tasks through software actions. In future research, this combination will allow the exploration of novel multi-agent LLM applications with advanced reasoning requirements and, overall, support systematic evaluations in cybersecurity and other domains.
Appendix A
Appendix A contains the result of generating an audit report by an agent of the type Audit Report Agent, which was previously invoked by an agent of the type Network Security Agent. The full results are published in the online repository at https://github.com/fhaer/multi-agent-llm-system.
[Uncaptioned image] Figure 3: Excerpt of a security report generated by an Audit Report Agent after the completion of server-side tasks by a Network Security Agent in the user interface of the client application.

### v4

@@ IV-B2 Reasoning Chains
Based on a series of prompts and responses, reasoning chains can be constructed in combination with constraints [16, 14]. In addition to Chain-of-Thought (CoT), which aims to issue prompts that generate subsequent reasoning steps with their responses, the chain can also be constructed explicitly by relying on pre-defined structures. Constraints ensure their existence and allow prompting the LLM of an agent with the next step. This approach is especially relevant in multi-agent scenarios, where another, specialized agent might be invoked based on data collected beforehand.
As an example, Listing 2 shows the Server-Security-Agent type that establishes an IP address, network, and scan result data after a series of prompts and invokes an agent of type Audit-Report-Agent. On the basis of these prompts, the agent is requested to create a report using the passed data. Further actions may impose constraints or specific conditions that influence which agent is invoked or the data passed to the agent. Figure 3 shows an excerpt of the generated report.
@@ V Conclusion
This paper presented the initial results of experimental research on the specification of multi-agent LLM systems. In this first system overview, the architecture and specification language were demonstrated and applied in test cases related to question answering as well as network and server security tasks. The test cases indicate the feasibility of the architecture and specification, showing the completion of tasks with agent specifications based on state-of-the-art commercially and openly available LLMs. In particular, the completion of the test cases shows the potential of LLM agents for software-based tasks in applications. Factors enabling the task completion are a combination of (1.) involving the knowledge capabilities of specialized agents such as for cybersecurity and code generation, (2.) utilizing reasoning techniques such as reasoning chains, and (3.) agents executing the inferred tasks through software actions. In future research, this combination will allow the exploration of novel multi-agent LLM applications with advanced reasoning requirements and, overall, support systematic evaluations in cybersecurity and other domains.
Figure 3: Excerpt of a security report generated by an Audit Report Agent after the completion of server-side tasks by a Network Security Agent in the user interface of the client application. The full results are published in the online repository at https://github.com/fhaer/multi-agent-llm-system.